### v2

# Quantifying Decoherence

Abstract
Quantum decoherence refers to the phenomenon when the interaction of a quantum system with its environment results in the degradation of quantum coherence. Decoherence is considered to be the most popular mechanism responsible for the emergence of classicality from quantum mechanics. The issue of formulating a measure of decoherence is addressed here. The approach taken here is that decoherence results from the entanglement of a quantum system with certain environment degrees of freedom, and quantifying this entanglement should yield the most natural measure of decoherence. A simple measure of decoherence is presented based on this notion, and it is examined for various example systems. The measure proves to be effective and is relatively straightforward to compute. In addition, a method has been proposed to measure decoherence in a Mach-Zehnder interferometer.
I Introduction
Emergence of classicality from quantum mechanics is an issue which has been debated right from its inception. One approach aimed at tackling this problem employs a form of non-linearity in quantum evolution, potentially leading to the dynamic evolution of macroscopic superposition states into a single, distinct macroscopic state [1, 2, 3]. The origin of this nonlinearity, however, is debatable. One view attributes it to an inherent non-linearity in the evolution equation [4]. Another considers it resulting from a gravitational self-interaction [5, 6, 7]. The behavior of a quantum system that is weakly interacting with numerous of degrees of freedom, referred to as the environment, has been extensively examined. Its suggested relationship with the emergence of classicality has given rise to a vibrant field known as decoherence [8, 9, 10]. The fundamental concept of the decoherence framework posits that classicality arises as a property of systems that interact with an environment, which effectively kills quantum coherence. Both qualitative and quantitative analyses of decoherence have yielded significant insights into the mechanisms behind the loss of quantum coherence, and several of its predictions have been validated through experimental testing [11, 12, 13]. Given that decoherence can disrupt fragile quantum characteristics and, consequently, the operation of devices that rely on quantum coherence for information processing, its investigation is of paramount importance for all experimental applications of quantum information and computation [13, 14, 15, 16].
Early works on decoherence mostly focused on the time scales over which decoherence occurs [9], which is interesting for addressing the question if the emergence of classicality is instantaneous or occurs over a time scale. With it also arose a need to quantify how classical a state becomes as it interacts with the environment. Lee [17] introduced a measure of “non-classicality” of a quantum state $\rho$, the quantity $\tau_{\rho}$, referred to as the “non-classical depth.” This measure was later used by several authors. The subject of quantifying decoherence has been previously reviewed [18, 19]. In more recent studies Zhang and Luo introduced an interesting decoherence measure in terms of the averaged Wigner-Yanase skew information [20], Fu and Luo approached quantifying decoherence via increase in classicality [21], Gundhi and Ulbricht introduced a quantifier named “decoherence kernel” [22]. It has been experimentally demonstrated that partially coherent electron wave packets can be constructed, which highlights the necessity to develop methods for quantifying decoherence [23].
In this work we consider decoherence as a phenomenon arising out of a quantum system getting entangled to certain environmental degrees of freedom through the weakest of interactions. We use this entanglement to quantify the degree of decoherence of the system.
II Decoherence in a nutshell
The process of decoherence can be generically represented in the following way. Let the state of a system and its environment be written as
$|\psi\rangle=\Big(\sum_{j=1}^{n}c_{j}|p_{j}\rangle\Big)|{\cal E}^{0}\rangle,$ (1)
where the state of the system is expanded in terms of certain basis states $|p_{j}\rangle$, and $|{\cal E}^{0}\rangle$ is the state of the environment. Assume a Hamiltonian and time evolution of the form .
${\mathscr{H}}$ $=$ $\sum_{j=1}^{n}|p_{j}\rangle\langle p_{j}|\otimes{\mathscr{H}}^{(j)}$ (2)
${\mathcal{U}}_{t}$ $=$ $\sum_{j=1}^{n}|p_{j}\rangle\langle p_{j}|\otimes{\mathcal{U}}_{t}^{(j)}$ (3)
where ${\mathcal{U}}_{t}^{(j)}=e^{-i{\mathscr{H}}^{(j)}t/\hbar}$ and the ${\mathscr{H}}^{(j)}$ are certain unspecified Hermitian operators involving the environment. It is obvious that ${\mathscr{H}}^{(j)}$ will depend on the choice of the basis states $|p_{j}\rangle$. The state (1), with the above specified time-evolution, evolves into
$|\psi(t)\rangle=\sum_{j=1}^{n}c_{j}|p_{j}\rangle{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle.$ (4)
The state (4) represents a grandiose entangled state involving the system and all of the environment. Practically it is impossible to keep track of all environmental degrees of freedom, and to probe them experimentally. That is the reason such a state is rarely used to describe decoherence. But this entanglement is essentially what leads to decoherence. This entangled state often leads to a many-worlds type of interpretation [24]. However, if one wants to look only at the system and forget about the environment, it is useful to write the reduced density matrix, which is obtained by writing the full density matrix for (4) as $\rho_{f}=|\psi(t)\rangle\langle\psi(t)|$ and tracing over the states of the environment, and has the form
$\rho(t)=\sum_{j,k}c_{k}^{*}c_{j}|p_{j}\rangle\langle p_{k}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle.$ (5)
Under the kind of time evolution specified above, the diagonal components of the density matrix, in the basis $|p_{j}\rangle$, remain unchanged, while off-diagonal elements are reduced by a factor $\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle\leq 1$. The temporal dependence of the suppressive factors will, in most cases, rely on the particular model of the environment and its interaction with the system. The idea of decoherence is that, for a specific basis $\{|p_{j}\rangle\}$, the suppressing terms rapidly decay over short time scales. Thus, over a time scale, called decoherence time-scale, the off diagonal terms in (5) disappear for all practical purposes, and one is left with an approximately diagonal density matrix of the system
$\rho\approx\sum_{j=1}^{n}|c_{j}|^{2}|p_{j}\rangle\langle p_{j}|.$ (6)
The off-diagonal elements related to quantum superpositions are now absent. Consequently, the system appears to exhibit classical behavior. This particular basis is chosen by the interaction with the environment, resulting in the emergence of classical states. In the literature, these states are referred to as pointer states, and the process of selection induced by the environment is termed einselection.
III Quantifying decoherence via entanglement
III.1 Entanglement between the system and environment
From the above discussion it is clear that decoherence is the result of the system getting entangled with the environment. Our approach is that the degree of decoherence can be quantified by quantifying the entanglement between the system and environment. In this combined system-environment aggregate, involving a lot of degrees of freedom, we consider a bipartition of the system-environment aggregate such that our system of interest forms one partition, and the rest of the environment forms the other. One can now ask, how much is the entanglement between the system and the rest of the environment. The generalized concurrence for this bipartition can be written as [25]
$E^{2}=4\sum_{j<k}(\rho_{jj}\rho_{kk}-\rho_{jk}\rho_{kj})=2[1-\text{tr}(\rho^{2})],$ (7)
where $\rho\equiv\sum_{i}\langle e_{i}|(|\psi\rangle\langle\psi|)|e_{i}\rangle=\text{Tr}_{e}(\rho_{f})$, $\{|e_{i}\rangle\}$ being an orthonormal basis for the whole of environment. This suggests that if the system has a finite dimensional Hilbert space, we can define a measure of decoherence as the normalized generalized concurrence for entanglement between the system and the environment. This decoherence measure is then given by
$\mathcal{D}_{e}\equiv\tfrac{n}{n-1}[1-\text{tr}(\rho^{2})],$ (8)
where $\rho$ is the reduced density operator of the system, after tracing over the environment states. In order to get the degree of decoherence of the system after a time $t$, in the scenario described above, we insert (5) in (8) to get
$\mathcal{D}_{e}$ $=$ $\tfrac{n}{n-1}[1-\sum_{i=1}^{n}\langle p_{i}|\sum_{j,k}c_{k}^{*}c_{j}|p_{j}\rangle\langle p_{k}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle$
$\sum_{l,m}c_{m}^{*}c_{l}|p_{l}\rangle\langle p_{m}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(m)}}^{\dagger}{\mathcal{U}}_{t}^{(l)}|{\cal E}^{0}\rangle|p_{i}\rangle$
$=$ $\tfrac{n}{n-1}\Big[1-\sum_{i,k}|c_{k}|^{2}|c_{i}|^{2}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(i)}|{\cal E}^{0}\rangle|^{2}\Big]$
$=$ $\tfrac{n}{n-1}\Big[1-\sum_{i=1}^{n}|c_{i}|^{4}-\sum_{j\neq k}|c_{j}|^{2}|c_{k}|^{2}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle|^{2}\Big].$
If the system states do not get correlated to the environment at all, it implies that $\mathcal{U}_{t}^{(k){\dagger}}\mathcal{U}_{t}^{(j)}=1$ for all $j,k$. In this situation $\mathcal{D}_{e}=\tfrac{n}{n-1}\big[1-\sum_{i=1}^{n}|c_{i}|^{4}-\sum_{j\neq k}|c_{j}|^{2}|c_{k}|^{2}\big]=0$. If the environment states correlated to the states $|p_{i}\rangle$ are all orthogonal to each other, and $|c_{i}|^{2}=\tfrac{1}{n}$ for all $i$, $\mathcal{D}_{e}=\tfrac{n}{n-1}[1-\sum_{i=1}^{n}|c_{i}|^{4}]=1$. This is the case of maximal decoherence of the system. Here the reduced density operator of the system is maximally mixed.
III.2 A qubit coupled to environment of spin-1/2
In the following we will use the measure introduced in the previous subsection to a single qubit coupled to a model environment of $N$ spin-1/2 entities. This system has already been studied in detail by Zurek [26]. Let the Hamiltonian be represented by
$H$ $=$ $\sum_{j}\epsilon_{j}|s_{j}\rangle\langle s_{j}|+\sum_{k}\mathcal{E}_{k}|e_{k}\rangle\langle e_{k}|$ (10)
$+\sum_{j,k}\gamma_{jk}|s_{j}\rangle\langle s_{j}|\otimes|e_{k}\rangle\langle e_{k}|,$
where $\epsilon_{j}$ is the energy eigenvalue of the j’th energy level of the system, $\mathcal{E}_{k}$ is the energy of the k’th state of the environment, $\gamma_{jk}$ the coupling strength between the respective states of the system and environment. Let the initial system-environment state be:
$|\Psi(0)\rangle$ $=$ $|\psi_{0}\rangle\otimes|\mathcal{E}_{0}\rangle$ (11)
$=$ $\sum_{j}\alpha_{j}|s_{j}\rangle\sum_{k}\beta_{k}|e_{k}\rangle.$
This state will evolve in time via the full Hamiltonian (10). The reduced density operator of the system, after a time $t$, is given by
$\rho(t)=\text{Tr}_{e}[|\Psi(t)\rangle\langle\Psi(t)|)=\sum_{i,j}\rho_{ij}(t)|s_{i}\rangle\langle s_{j}|,$ (12)
where
$\rho_{ii}(t)$ $=$ $|\alpha_{i}|^{2}\sum_{k}|\beta_{k}|^{2}=|\alpha_{i}|^{2}$
$\rho_{ij}(t)$ $=$ $\alpha_{i}\alpha_{j}^{*}e^{-it(\epsilon_{i}-\epsilon_{j})}\sum_{k}|\beta_{k}|^{2}e^{-it(\gamma_{ik}-\gamma_{jk})}.$ (13)
One can now evaluate the following useful term
$\text{tr}(\rho^{2})$ $=$ $\sum_{i,j}|\alpha_{i}|^{2}|\alpha_{j}|^{2}\times$ (14)
$\sum_{k}|\beta_{k}|^{2}e^{-it(\gamma_{ik}-\gamma_{jk})}\sum_{l}|\beta_{l}|^{2}e^{it(\gamma_{il}-\gamma_{jl})}.$
If one define one of the sums in the above expression by
$z_{ij}=\sum_{k}|\beta_{k}|^{2}e^{-it(\gamma_{ik}-\gamma_{jk})},$ (15)
It has been shown that the time average of it, over a reasonably long span of time $T$, is [26]
$\langle z_{ij}\rangle_{T}\to 0,$
$\langle|\langle z_{ij}|^{2}\rangle_{T}\to\frac{1}{N}.$ (16)
The decoherence measure for the system, at a time $t$, can now be evaluated as
$\mathcal{D}_{e}$ $=$ $\tfrac{n}{n-1}[1-\text{tr}(\rho^{2})]$ (17)
$=$ $\tfrac{n}{n-1}\Big[1-\sum_{i}|\alpha_{i}|^{4}-\sum_{i\neq j}|\alpha_{i}|^{2}|\alpha_{j}|^{2}\langle|\langle z_{ij}|^{2}\rangle_{T}\Big]$
$\approx$ $\tfrac{n}{n-1}\Big[1-\sum_{i}|\alpha_{i}|^{4}-\frac{1}{N}\sum_{i\neq j}|\alpha_{i}|^{2}|\alpha_{j}|^{2}\Big].$
If the environment is large, the last term in the above is negligible, and the decoherence measure becomes
$\mathcal{D}_{e}$ $\approx$ $\tfrac{n}{n-1}[1-\sum_{i}|\alpha_{i}|^{4}].$ (18)
This is the maximum value the decoherence measure will attain, for a given $\{\alpha_{i}\}$. If all $|\alpha_{i}|^{2}=1/n$, the decoherence measure will attain its maximal value $\mathcal{D}_{e}\approx 1$.
III.3 Decoherence in the spin-boson model
Next we analyze a model, studied very well in the context of quantum dissipation, namely the spin-boson model [27]. It involves a two-level system linearly coupled a to bath of harmonic oscillators, so that the total Hamiltonian is
$\mathcal{H}$ $=$ $\tfrac{\omega_{s}}{2}\sigma_{z}+\sum_{k}\omega_{k}b_{k}^{{\dagger}}b_{k}+\sum_{k}\sigma_{z}(g_{k}b^{{\dagger}}_{k}+g_{k}^{*}b_{k}),$ (19)
where $\sigma$ represents an operator for the two-level system, $b_{k}$ are the operators for the harmonic oscillators, and $g_{k}$ are certain coupling constants. This model was introduced to study the quantum dissipative dynamics of a two-level system. However, here we are just concerned with the decoherence caused by the bath of harmonic oscillators. The effect of the environment on the two-level system can studied by analyzing the reduced density operator obtained after tracing over the bath states. The reduced density operator follows the equation: [27]
$\frac{\partial\rho(t)}{\partial t}$ $=$ $-\tfrac{i\omega_{s}}{2}[\sigma_{z},\rho(t)]+\gamma\sigma_{z}\rho(t)\sigma_{z}-\gamma\rho(t),$ (20)
where $\gamma$ contains all the effect of the heat-bath on the system. There is a well established procedure to obtain $\gamma$ by assuming a kind of heat-bath energies and the couplings $g_{k}$ [27]. For an initial equal superposition of the eigenstates of $\sigma_{z}$, the reduced density operator of the system can be obtained as
$\rho(t)$ $=$ $\frac{1}{2}\begin{pmatrix}1&e^{-i\omega_{s}t}e^{-2\gamma t}\\ e^{i\omega_{s}t}e^{-2\gamma t}&1\\ \end{pmatrix}$. (21)
The decoherence measure can be evaluated from the above to yield
$\mathcal{D}_{e}$ $=$ $1-e^{-4\gamma t}.$ (22)
We note that as time progresses, the decoherence measure $\mathcal{D}_{e}$ approaches its maximal value 1.
IV Quantifying decoherence for continuous variable systems
IV.1 Generalized entanglement measure and decoherence
For continuous variable systems, we follow the same philosophy that we followed while quantifying decoherence for finite dimensional systems. We treat the system-environment combine as an entangled system, and divide into two partitions. The system of interest constitutes one partition, and the environment constitutes the other. We quantify the entanglement across this bipartition using the generalized entanglement measure (GEM) introduced by Swain, Bhaskara, and Panigrahi [28]. Consider a $n$-partite entangled state
$|\psi\rangle=\int\psi(x_{1},x_{2},\dots,x_{n})|x_{1}\rangle|x_{2}\rangle\dots|x_{n}\rangle dx_{1}dx_{2}\dots dx_{n},$
with $\int\psi^{*}(x_{1},x_{2}\dots,x_{n})\psi(x_{1},x_{2}\dots,x_{n})d^{n}x=1$. Consider a bipartition $\mathcal{M}$ of the combined Hilbert space, one containing $m$ degrees of freedom, and the other containing nmn-m degrees of freedom. The GEM is defined for this $n$-partite entangled state, for the bipartition $\mathcal{M}$, as [28]
$\mathcal{E}_{\mathcal{M}}^{2}$ $=$ $2\Big[1-\int\Big|\int\psi(x_{1}^{\prime},\dots,x_{m}^{\prime},y_{m+1},\dots y_{n})$
$\psi^{*}(x_{1},\dots,x_{m},y_{m+1},\dots y_{n})dy_{m+1}\dots dy_{n}\Big|^{2}d^{m}xd^{m}x^{\prime}\Big],$
The GEM is bounded by 0 and 2. For our case of a single system interacting with a large environment of $n$ degrees of freedom, $m=1$. Since we prefer a normalized entanglement measure, we drop the 2 in the above expression, and define the decoherence measure as
$\mathcal{D}_{e}$ $\equiv$ $\Big[1-\int\Big|\int\psi(x^{\prime},y_{1},\dots y_{n})$ (25)
$\psi^{*}(x,y_{1},\dots y_{n})dy_{1}\dots dy_{n}\Big|^{2}dxdx^{\prime}\Big],$
$=$ $\Big[1-\int\big|\rho(x^{\prime},x)\big|^{2}dxdx^{\prime}\Big],$
where $\rho(x^{\prime},x)=\int\psi(x^{\prime},y_{1},\dots y_{n})\psi^{*}(x,y_{1},\dots y_{n})d^{n}y$ is the reduced density matrix for the system of interest after tracing over the environment states. In order to evaluate the decoherence measure, we just need to obtain the reduced density matrix. Eq. (25) can be written more compactly as $\mathcal{D}_{e}=1-\text{Tr}[\rho^{2}]$ which generalizes it to situations where the system of interest may involve both continuous and discrete degrees of freedom.
For continuous variable systems, figuring out the pointer basis is a difficult problem. Various studies have been done on model systems trying to find out the pointer states [26, 29, 30, 31, 32, 33, 34, 35, 36]. The decoherence measure defined by (25) offers the benefit that the knowledge of the pointer states is not required for its evaluation.
IV.2 Decoherence of a free particle
In order to apply the decoherence measure we a consider the simplest system with continuous variables. Specifically we consider a particle of mass $m$, traveling along x-axis. For simply modeling interaction with an environment, the particle is assumed to be interacting with bath of unit-mass harmonic oscillators by a coupling of its position to the momenta of the harmonic oscillators. If we just look at the motion of the particle in the x-direction, the Hamiltonian can be written as [37]
Figure 1: The decoherence measure $\mathcal{D}_{e}$ given by (29), plotted against $t/\tau_{D}$ for a free particle undergoing decoherence, which is initially in a plane wave state.
$H=\frac{P_{x}^{2}}{2m}+\sum_{k}(p_{k}^{2}+\tfrac{1}{2}\omega_{k}r_{k}^{2})+x\sum_{k}g_{k}p_{k},$ (26)
where $x,P_{x}$ represent the position and momentum operator of the particle, $r_{k},p_{k}$ are position and momentum of the k’th oscillator. This model was originally proposed for studying quantum dissipation or damping [37]. However, here we use it to study decoherence. The problem of a particle coupled to a a bath of independent harmonic oscillators was studied in detail by Caldeira and Leggett [38]. One can write a density operator for the particle and the heat-bath combined, and evolves it using the Hamiltonian (26). The dynamics of the particles can then be studied by averaging over the environment degrees of freedom, and looking at the reduced density operator of the particle. Under certain conditions the reduced density matrix of the particle follows the equation [38]
$\frac{\partial\rho}{\partial t}$ $=$ $\frac{i}{2m\hbar}[P_{x}^{2},\rho]-\frac{i\gamma}{\hbar}[x,P_{x}\rho+\rho P_{x}]-\frac{4\pi\gamma}{\lambda_{T}^{2}}[x,[x,\rho]],$
where $\gamma$ is a damping parameter similar to the one introduced in the preceding subsection, and $\lambda_{T}=\sqrt{\frac{2\pi\hbar^{2}}{mk_{B}T}}$ is the thermal deBroglie wavelength of the particle. We consider the particle initially to be in a plane wave state $\psi(x,0)=e^{ikx}$ with a wave-vector $k$, but confined to a spatially extended region from L-L to +L+L. The initial density matrix of the particle is $\frac{1}{2L}e^{ik(x-x^{\prime})}$. With this initial condition the equation (LABEL:master-sw) can be solved and the reduced density matrix of the particle at a later time $t$ is given by [37]
$\rho(x,x^{\prime},t)=\exp\{-ie^{-2\gamma t}k(x-x^{\prime})\}e^{-\sigma(x-x^{\prime})^{2}},$ (28)
where $\sigma=\frac{\pi}{\lambda_{T}^{2}}(1-e^{-4\gamma t})$. Before proceeding further we would like to emphasize the time scales one should look at while studying decoherence. In the problem studied by Savage and Walls [37], the parameter $\gamma$ has a simple meaning - it is the thermal relaxation rate of the particle due to the damping effect of the environment. However, while studying decoherence one is studying such minimal effects of the environment that no change in the energy of the particle results from it, and particle travels virtually undamped [36]. The effect of the environment is only a dephasing of its time evolution. Decoherence effects should come into play at times much shorter than $1/\gamma$ [36]. For $\gamma t\ll 1$ one can make the following approximation: $\sigma\approx\frac{4\pi}{\lambda_{T}^{2}}\gamma t$. One can now use (25) to evaluate the decoherence measure, and it comes out to be
$\mathcal{D}_{e}=1-\frac{\sqrt{\pi}}{2\sqrt{2\sigma}L}\text{erf}(\sqrt{2\sigma}L).$ (29)
Notice that in the limit $\sqrt{\sigma}L\gg 1$, i.e., for $t\gg\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ the decoherence measure $\mathcal{D}_{e}$ tends to its maximum value 1. The time scale that emerges here agrees with the decoherence time scale $\tau_{D}=\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ estimated earlier [9]. For intermediate times one can estimate the degree of decoherence, the particle has undergone, from (29). The decoherence measure is plotted as a function of time (scaled by $\tau_{D}$) in Fig. 1. It is interesting to see that the growth of the degree of decoherence with time is distinctly slower than the naive expectation of $1-e^{-t/\tau_{D}}$.
If one starts with a particle which is localized, its wavefunction may be described by a Gaussian wavepacket
$\psi(x,0)=\tfrac{1}{(\pi\sigma^{2})^{1/4}}e^{-x^{2}/2\sigma^{2}},$ (30)
where $\sigma$ is a parameter such that the position spread of the wavefunction is $2\sigma$. The effect of decoherence on the particle can be studied by solving (LABEL:master-sw) using the above initial state. The reduced density matrix of the particle at a time $t$, in the limit $\gamma t\ll 1$, is then given by [37]
$\rho(x,x^{\prime},t)$ $=$ $\tfrac{1}{\sqrt{\pi c}}e^{-ibrR/c},e^{-R^{2}/4c}e^{-(a-b^{2}/c)r^{2}}$ (31)
where $R=x+x^{\prime},~r=x-x^{\prime}$, and $a=1/4\sigma^{2}$, $b=\hbar t/2m\sigma^{2}$, $c=\sigma^{2}+(\frac{\hbar t}{\sigma m})^{2}+\frac{16k_{B}T\gamma t^{3}}{m}$. The decoherence measure can now be evaluated by using (25), and is given by
$\mathcal{D}_{e}=1-\frac{1}{\sqrt{1+\frac{16k_{B}T\gamma t^{3}}{m\sigma^{2}}}}.$ (32)
In the limit $t\to 0$, the decoherence measure $\mathcal{D}_{e}=1$. For $t\gg(\frac{m\sigma^{2}}{16k_{B}T\gamma})^{1/3}$ the decoherence measure approaches its maximal value 1. Interestingly, here the decoherence is not faster if the position spread of the particle is large, i.e., $\sigma$ is large. Rather, it is faster if the momentum spread of the particle is large, i.e., $\hbar/\sigma$ is large. This is probably an artifact of the coupling of the particle’s position to the momenta of the environment, in (26).
Figure 2: The decoherence measure $\mathcal{D}_{e}$ given by (36), plotted against $t/\tau_{D}$ for a free particle undergoing decoherence, which is initially in a plane wave state.
A more realistic coupling to the environment is given by the Hamiltonian [38]
$H=\frac{P_{x}^{2}}{2m}+\sum_{k}\left(\tfrac{1}{2}m_{k}\omega_{k}^{2}(r_{k}-\tfrac{g_{k}x}{m_{k}\omega_{k}})^{2}+\frac{p_{k}^{2}}{2m_{k}}\right).$ (33)
The time evolution of the reduced density matrix is governed by the following master equation [38]
$\frac{\partial\rho(x,x^{\prime},t)}{\partial t}$ $=$ $\Bigg\{\frac{-\hbar}{2im}\left(\frac{\partial^{2}}{\partial x^{2}}-\frac{\partial^{2}}{\partial x^{\prime 2}}\right)-\gamma(x-x^{\prime})$
$\left(\frac{\partial}{\partial x}-\frac{\partial}{\partial x^{\prime}}\right)-\tfrac{D}{4\hbar^{2}}(x-x^{\prime})^{2}\Bigg\}\rho(x,x^{\prime},t),$
where the parameters are the same as described in the preceding analysis. Starting from an initial Gaussian state (30), the reduced density matrix of the particle, in the limit $\gamma t\ll 1$, is given by [36]
$\rho(R,r,t)$ $=$ $\tfrac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\tfrac{2m\gamma k_{B}Tr^{2}}{\hbar^{2}}t-\tfrac{R^{2}+r^{2}}{4\sigma^{2}}\right)$
The decoherence measure, evaluated by using (25), is given by
$\mathcal{D}_{e}=1-\frac{1}{\sqrt{1+t/\tau_{D}}},$ (36)
where
$\tau_{D}=\frac{\hbar^{2}}{2m\gamma k_{B}T4\sigma^{2}}.$ (37)
This is consistent with the existing form of the decoherence time-scale, as the wave-packet is $2\sigma$ wide. However, the increase in decoherence with time is far from exponential.
IV.3 Decoherence in a Stern-Gerlach experiment
Next we look at an experiment which is a prototype for studying quantum measurement, namely, the Stern-Gerlach experiment. In this experiment we consider a spin-1/2 particle of mass $m$, traveling along y-axis, and passing through an inhomogeneous magnetic field pointing along z-axis. The field is inhomogeneous in the x-direction. Without going into calculational details, we use this setup to illustrate the usefulness the present decoherence measure in dealing with situations where there both discrete and continuous degrees of freedom. Two different kinds of decoherence effects are expected to happen in this experiment, and at two different time scales. It is interesting to see how the decoherence measure captures the two effects. If we just look at the evolution of the particle in the x-direction, the Hamiltonian can be written as [30]
$H=\frac{p_{x}^{2}}{2m}+\lambda\sigma_{z}+\epsilon x\sigma_{z}+H_{SE}+H_{E},$ (38)
where $x,p_{x}$ represent the position and momentum operator of the particle, $\epsilon$ the product of the field gradient and the magnetic moment of the particle. The environment is assumed to be a bath of independent harmonic oscillators, with the Hamiltonian $H_{E}$. The particle interacts with this environment through an interaction term $H_{SE}$. One writes a density operator for the particle and the heat-bath combined, and evolves it using the Hamiltonian (38). The dynamics of the particles can then be studied by averaging over the environment degrees of freedom, and looking at the reduced density operator of the particle. The reduced density matrix of the particle follows the equation [30]
$\frac{\partial\rho_{ss^{\prime}}(x,x^{\prime},t)}{\partial t}$ $=$ $\Bigg\{\frac{-\hbar}{2im}\left(\frac{\partial^{2}}{\partial x^{2}}-\frac{\partial^{2}}{\partial x^{\prime 2}}\right)-\gamma(x-x^{\prime})$
$\left(\frac{\partial}{\partial x}-\frac{\partial}{\partial x^{\prime}}\right)-\tfrac{\pi\gamma}{\lambda_{T}^{2}}(x-x^{\prime})^{2}+$
$\frac{i\epsilon(xs-x^{\prime}s^{\prime})}{\hbar}+\frac{i\lambda(s-s^{\prime})}{\hbar}\Bigg\}\rho_{ss^{\prime}}(x,x^{\prime},t),$
where $s,s^{\prime}$ are the spin labels $\pm$, $\gamma$ and $\lambda_{T}$ are the relaxation rate and the thermal deBroglie wavelength, as discussed in (LABEL:master-sw). The effect of the inhomogeneous magnetic field that the particle described by a localized Gaussian wave-packet evolves into a superposition of two separated wave-packets, which are correlated the spin states $|\pm\rangle$ defined by $\sigma_{z}|\pm\rangle=\pm|\pm\rangle$. In the spin space, the reduced density operator, at a later time $t$, can be written as
$\rho(t)$ $=$ $\begin{pmatrix}\rho_{++}&\rho_{+-}\\ \rho_{-+}&\rho_{--}\\ \end{pmatrix}$. (40)
It has been argued earlier [9] that the decoherence time-scale $\tau_{D}$ is inversely proportional to the spatial separation between the two points between which one is looking at decoherence. In the spin-diagonal components of the density matrix $\rho_{++},\rho_{--}$, the separation between $x$ and $x^{\prime}$ is of the order of the width of a Gaussian wave-packet. On the other hand, in off-diagonal components of the density matrix $\rho_{+-},\rho_{-+}$, the separation between $x$ and $x^{\prime}$ is of the order of the separation between the two spatially separated Gaussians. Thus $\rho_{+-},\rho_{-+}$ decay at a much shorter time scale. Eventually $\rho_{++},\rho_{--}$ also experience decoherence, but over a relatively longer time-scale.
The decoherence measure is given by $\mathcal{D}_{e}=1-\text{Tr}[\rho^{2}]$, where the trace now involves a trace over the spin state and a trace over the spatial degrees of freedom. Partial trace over the spin states of $\rho^{2}$ gives
$\text{Tr}_{s}[\rho(t)^{2}]=\rho_{++}^{2}+\rho_{--}^{2}+\rho_{+-}\rho_{-+}+\rho_{-+}\rho_{+-}.$ (41)
The full trace of $\rho^{2}$ can then be performed, and the decoherence measure can be evaluated as
${D}_{e}$ $=$ $1-\text{Tr}[\rho(t)^{2}]$ (42)
$=$ $1-\int(|\rho_{++}(x,x^{\prime})|^{2}+|\rho_{--}(x,x^{\prime})|^{2})dxdx^{\prime}$
$-2\int|\rho_{+-}(x,x^{\prime})|^{2}dxdx^{\prime}$
One can see that over a short time scale the last term in the above will decay to a negligible value, and the decoherence measure will increase, but will not reach its maximal value. Over a longer time-scale the terms involving $\rho_{++},\rho_{--}$ will also experience a decay, as seen in the case of decoherence of a Gaussian wavefunction in the preceding subsection, and the decoherence measure will approach its maximal value. The decoherence measure is thus well suited to quantify decoherence setting in progressively over different time scales.
V Measuring decoherence
Figure 3: A schematic diagram of a Mach-Zehnder interferometer. The quanton, as it traverses the two paths, may undergo decoherence as a result of interaction with the environment.
In the following we propose a method of experimentally measuring decoherence in a Mach-Zehnder interferometer. Consider a Mach-Zehnder interferometer as shown in Fig. 3, where the quantons that travel through it may be photons, neutrons or any other microparticle. The quanton first encounters the beam-splitter BS1 and is split into a superposition of two paths. Let us assume that the state of the quanton, after passing through BS1, is
$|\psi_{0}\rangle=\tfrac{1}{\sqrt{2}}(|\psi_{1}\rangle+|\psi_{2}\rangle)$ (43)
By the time the quanton reaches the beam-split BS2, the phase difference between the paths may have changed. We assume an additional phase difference between the two paths so that the state is now
$|\psi\rangle=\tfrac{1}{\sqrt{2}}(|\psi_{1}\rangle+e^{i\phi}|\psi_{2}\rangle).$ (44)
The effect of BS2 on the quanton is such that it splits a path into a superposition of two paths, one traveling to the detector $D1$ and the other traveling to the detector $D2$. However there is a difference in the relative phases of the split paths depending on whether the quanton came from the upper path (1) or the lower path (2). The action of BS2 can be summarized as follows
$U_{BS2}|\psi_{1}\rangle$ $=$ $\tfrac{1}{\sqrt{2}}(|D_{1}\rangle+|D_{2}\rangle)$
$U_{BS2}|\psi_{2}\rangle$ $=$ $\tfrac{1}{\sqrt{2}}(|D_{1}\rangle-|D_{2}\rangle).$ (45)
In optics this happens when the lower surface of BS2 is silvered. So after passing through BS2, the quanton will be in the state
$|\psi_{f}\rangle$ $=$ $U_{BS2}\tfrac{1}{\sqrt{2}}(|\psi_{1}\rangle+e^{i\phi}|\psi_{2}\rangle)$ (46)
$=$ $\tfrac{1}{2}(|D_{1}\rangle+|D_{2}\rangle)+e^{i\phi}\tfrac{1}{2}(|D_{1}\rangle-|D_{2}\rangle).$
The probability of the quanton hitting D1 is $(1+\cos\phi)/2$, and that of hitting D2 is $(1-\cos\phi)/2$. If $\phi=0$ the probability of the quanton hitting D2 is zero, and all the quantons will hit D1. This represents sharp interference. Small nonzero values of $\phi$ lead to unsharp interference.
Now if the quanton, while traversing the two paths, interacts with the environment, and undergoes decoherence by the time it reaches BS2, it cannot be represented by a pure state. One can trace over the environment states, and obtain a reduced density matrix for the quanton. One can look at the density matrix in the 2-dimensional Hilbert space spanned by $|\psi_{1}\rangle,|\psi_{2}\rangle$. In this space it looks like the following:
$\rho$ $=$ $\begin{pmatrix}\rho_{11}&\rho_{12}e^{-i\phi}\\ \rho_{21}e^{i\phi}&\rho_{22}\\ \end{pmatrix}$, (47)
where $\rho_{11}$ ($\rho_{22}$) represents the probability of the quanton passing through the upper (lower) path. The decoherence measure can then be calculated using (8) as
${D}_{e}$ $=$ $2(1-\rho_{11}^{2}-\rho_{22}^{2}-2|\rho_{12}|^{2}).$ (48)
The problem at hand now is find a way to make measurements in the experiment to evaluate the above decoherence measure. If one blocks the lower path before it enters BS2, there will be equal intensity at D1 and D2 which will be proportional to $\rho_{11}$. On the other hand, if one blocks the upper path before it enters BS2, there will again be equal intensity at D1 and D2 which will now be proportional to $\rho_{22}$. It can be shown that the average intensity at the two detectors, in the two scenarios, is given by
$(I_{av})_{1}=\tfrac{1}{2}(I_{D1}+I_{D2})_{1}$ $=$ $\rho_{11}$
$(I_{av})_{2}=\tfrac{1}{2}(I_{D1}+I_{D2})_{2}$ $=$ $\rho_{22},$ (49)
where $I_{D1},I_{D2}$ represent the intensities at the two detectors, and the subscripts after the brackets denote which path is open. When both the paths are open, the intensities at the two detectors are given by $I_{D1}=\rho_{11}+\rho_{22}+\rho_{12}e^{-i\phi}+\rho_{21}e^{i\phi}$ and $I_{D2}=\rho_{11}+\rho_{22}-\rho_{12}e^{-i\phi}-\rho_{21}e^{i\phi}$. The difference in the intensities at the two detectors is then given by
$I_{diff}=\tfrac{1}{2}(I_{D1}-I_{D2})=2|\rho_{12}|\cos\phi$ (50)
The decoherence measure can now be evaluated as
${D}_{e}$ $=$ $2[1-(I_{av})_{1}^{2}-(I_{av})_{2}^{2}-\tfrac{1}{2}I_{diff}^{2}],$ (51)
where we have assumed $\phi=0$. If one is unsure about $\phi$ in the setup, one can tune the phase shifter such that maximum value of $I_{diff}^{2}$ is obtained. In this analysis it is assumed that all intensities are normalized with the total intensity. Given that the Mach-Zehnder interferometer with cold neutrons is widely used and there have been significant developments in the field of neutron interferometry [39], the experiment suggested here should be feasible with neutrons, allowing for the measurement of decoherence. The investigation of decoherence through neutron interferometry has been previously suggested [40]. While decoherence in a conventional photonic Mach-Zehnder interferometer in a lab would be negligible, one may use the proposed method to measure decoherence in optical interferometry involving large distances, e.g., as in the setup involving gravitational wave detection. Interferometry with atoms and molecules is a topic of active research, and decoherence is an important issue there [41, 42, 43, 44]. One can think of extending the method proposed here to such matter-wave interferometry experiments.
VI Discussion and conclusions
To summarize, we formulated a measure of decoherence based on the quantification of entanglement of a system with the environment. The measure is given by $\mathcal{D}_{e}=\tfrac{n}{n-1}[1-\text{tr}(\rho^{2})]$ for systems with finite-dimensional Hilbert space, and $\mathcal{D}_{e}=[1-\text{tr}(\rho^{2})]$ for continuous variable systems, where $\rho$ represents the reduced density operator of the system after tracing over the environment degrees of freedom. This decoherence measure has minimum value 0 and maximum value 1. Figuring out the pointer states that emerge out of decoherence of a system is often a difficult issue. In one basis, the off-diagonal elements might diminish more rapidly over time, whereas in another basis, they may also diminish, albeit at a slower rate [36]. The present formulation does not depend on the pointer states. One just needs to evaluate a trace which is basis independent.
Some earlier works treat the classicality of a state as quantifying the degree of decoherence. In our view, these two are separate things. A state may be close to a classical state, e.g., a coherent state for a harmonic oscillator, but it may not have arisen via decoherence. The classicality of such a state may not be a measure of decoherence.
We have also proposed a method of measuring decoherence in a Mach-Zehnder interferometer. The method is straightforward, and just involves measuring intensities in a few different settings. We believe the measure of decoherence introduced here will turn out to be useful in quantifying decoherence in different situations, especially in the quantum information related experiments.
Acknowledgements.
MSQ acknowledges financial support from the University Grants Commission, India, through the Junior Research Fellowship (NTA reference number: 231610153280).

### v3

@@
 The measure proves to be effective and is relatively straightforward to compute.
-In addition, a method has been proposed to measure decoherence in a Mach-Zehnder interferometer.
+In addition, a method has been proposed to measure decoherence in a Mach-Zehnder interferometer which may be useful in neutron interferometry.
 I Introduction
@@
 It has been experimentally demonstrated that partially coherent electron wave packets can be constructed, which highlights the necessity to develop methods for quantifying decoherence [23].
+In an interesting study, decoherence induced during the preparation of highly non-classical states for quantum metrology, arising from entanglement with a quantized auxiliary system was studied using quantum Fisher information (QFI) [24].
+This approach applies to situations where the auxillary system cannot be probed, which is also the case with the environment in standard decoherence.
+However it seems more suited to specific scenarios like state preparation, and may not apply to, say, environment induced decoherence of a particle passing through a double slit.
+Nevertheless, QFI has become a popular tool to probe decoherence in many subsequent studies.
+About a decade back, a specific measure of coherence was established, primarily driven by its necessity in the domain of quantum information [25].
+This measure has demonstrated significant utility, even in quantifying the wave characteristics of quantum particles [26].
+One might imagine that, given the strong association between decoherence and the loss of coherence, one could express decoherence in terms of the coherence measure.
+Nevertheless, the challenge in this approach lies in the fact that the measure of coherence is dependent on the basis, while it is desirable to have a measure of decoherence that is independent of the basis, for reasons that will become clear in the subsequent discussion.
+As an open quantum system evolves, it typically establishes correlations with its environment and irreversibly diminishes its information content.
+Nevertheless, this holds true only when the system is weakly coupled to the environment.
+If the coupling of the system to the environment is strong, information may periodically revert to its original state from the environment, resulting in a back-flow of information.
+This phenomenon is referred to as non-Markovian behavior.
+Of late quantum non-Markovianity has become an active subject of research [27, 28, 29], and measures of the degree of non-Markovianity have been proposed [30, 31].
+However, the decoherence induced emergence of classicality arises from an extremely weak interaction with the environment, so minimal that the system effectively evolves without dissipation.
+Consequently, non-Markovianity may not be applicable in these contexts.
+Therefore, it is necessary to establish a measure that is relevant to scenarios involving weak interactions with the environment.
 In this work we consider decoherence as a phenomenon arising out of a quantum system getting entangled to certain environmental degrees of freedom through the weakest of interactions.
@@
 One can now ask, how much is the entanglement between the system and the rest of the environment.
-The generalized concurrence for this bipartition can be written as [25]
+To this end we use the methodology developed by Bhaskara and Panigrahi [33] for quantifying multipartite entanglement.
+Let us assume that our system of interest has interacted with certain environment degrees of freedom, such that the combined state of the system and the environment is given by (4).
+We assume the the degrees of freedom described by the states $|p_{i}\rangle$ form one partition, and those described by the states $\mathcal{U}_{t}^{(i)}|\mathcal{E}^{0}\rangle$ form the other much larger partition.
+We wish to quantify the entanglement between these two partitions.
+In the methodology of Bhaskara and Panigrahi [33], one first traces over the degrees of freedom of one partition to obtain a reduced density operator.
+In our case, we trace over the environment states and get the reduced density operator $\rho$ for our system of interest.
+The generalized concurrence for this bipartition can then be written as [33]
 $E^{2}=4\sum_{j<k}(\rho_{jj}\rho_{kk}-\rho_{jk}\rho_{kj})=2[1-\text{tr}(\rho^{2})],$ (7)
@@
 where $\rho$ is the reduced density operator of the system, after tracing over the environment states.
-In order to get the degree of decoherence of the system after a time $t$, in the scenario described above, we insert (5) in (8) to get
+We first test out this measure for the scenario described above.
+In order to get the degree of decoherence of the system after a time $t$, we plug in the reduced density operator from (5) in the expression for the decoherence measure (8).
+That procedure results in following form of the decoherence measure:
 $\mathcal{D}_{e}$ $=$ $\tfrac{n}{n-1}[1-\sum_{i=1}^{n}\langle p_{i}|\sum_{j,k}c_{k}^{*}c_{j}|p_{j}\rangle\langle p_{k}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle$
@@
 $=$ $\tfrac{n}{n-1}\Big[1-\sum_{i=1}^{n}|c_{i}|^{4}-\sum_{j\neq k}|c_{j}|^{2}|c_{k}|^{2}|\langle{\cal E}^{0}|{{\mathcal{U}}_{t}^{(k)}}^{\dagger}{\mathcal{U}}_{t}^{(j)}|{\cal E}^{0}\rangle|^{2}\Big].$
-If the system states do not get correlated to the environment at all, it implies that $\mathcal{U}_{t}^{(k){\dagger}}\mathcal{U}_{t}^{(j)}=1$ for all $j,k$.
+If the system states do not get correlated to the environment at all, it implies that the environment states correlated to the system states, $\mathcal{U}_{t}^{(i)}|\mathcal{E}^{0}\rangle$, are all identical, and $\mathcal{U}_{t}^{(k){\dagger}}\mathcal{U}_{t}^{(j)}=1$ for all $j,k$.
 In this situation $\mathcal{D}_{e}=\tfrac{n}{n-1}\big[1-\sum_{i=1}^{n}|c_{i}|^{4}-\sum_{j\neq k}|c_{j}|^{2}|c_{k}|^{2}\big]=0$.
@@
 However, here we use it to study decoherence.
-The problem of a particle coupled to a a bath of independent harmonic oscillators was studied in detail by Caldeira and Leggett [38].
-One can write a density operator for the particle and the heat-bath combined, and evolves it using the Hamiltonian (26).
+The problem of a particle coupled to a bath of independent harmonic oscillators was studied in detail by Caldeira and Leggett [47, 48].
+One can write a density operator for the particle and the heat-bath combined, and evolve it using the Hamiltonian (26).
 The dynamics of the particles can then be studied by averaging over the environment degrees of freedom, and looking at the reduced density operator of the particle.
-Under certain conditions the reduced density matrix of the particle follows the equation [38]
+Under certain conditions the reduced density matrix of the particle follows the equation [47, 48]
 $\frac{\partial\rho}{\partial t}$ $=$ $\frac{i}{2m\hbar}[P_{x}^{2},\rho]-\frac{i\gamma}{\hbar}[x,P_{x}\rho+\rho P_{x}]-\frac{4\pi\gamma}{\lambda_{T}^{2}}[x,[x,\rho]],$
@@
 In the problem studied by Savage and Walls [46], the parameter $\gamma$ has a simple meaning - it is the thermal relaxation rate of the particle due to the damping effect of the environment.
-However, while studying decoherence one is studying such minimal effects of the environment that no change in the energy of the particle results from it, and particle travels virtually undamped [36].
+However, while studying decoherence one is studying such minimal effects of the environment that no change in the energy of the particle results from it, and particle travels virtually undamped [44, 45].
 The effect of the environment is only a dephasing of its time evolution.
@@
 Notice that in the limit $\sqrt{\sigma}L\gg 1$, i.e., for $t\gg\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ the decoherence measure $\mathcal{D}_{e}$ tends to its maximum value 1.
+Here $1/\sqrt{\sigma}$ is related to the width of the wavepacket at time $t$.
+So the limit implies the situation where decoherence is strong, and the wavepacket doesn’t spread much even after a time $t$.
 The time scale that emerges here agrees with the decoherence time scale $\tau_{D}=\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ estimated earlier [9].
@@
 $\mathcal{D}_{e}=1-\frac{1}{\sqrt{1+\frac{16k_{B}T\gamma t^{3}}{m\sigma^{2}}}}.$ (32)
-In the limit $t\to 0$, the decoherence measure $\mathcal{D}_{e}=1$.
-For $t\gg(\frac{m\sigma^{2}}{16k_{B}T\gamma})^{1/3}$ the decoherence measure approaches its maximal value 1.
+In the limit $t\to 0$, the decoherence measure $\mathcal{D}_{e}=0$, implying that the environment influence is negligible during such short times.
+For longer times the environment has enough time to cause decoherence, so for $t\gg(\frac{m\sigma^{2}}{16k_{B}T\gamma})^{1/3}$ the decoherence measure approaches its maximal value 1.
 Interestingly, here the decoherence is not faster if the position spread of the particle is large, i.e., $\sigma$ is large.
@@
 Figure 2: The decoherence measure $\mathcal{D}_{e}$ given by (36), plotted against $t/\tau_{D}$ for a free particle undergoing decoherence, which is initially in a plane wave state.
-A more realistic coupling to the environment is given by the Hamiltonian [38]
+A more realistic coupling to the environment is given by the Hamiltonian [47, 48]
 $H=\frac{P_{x}^{2}}{2m}+\sum_{k}\left(\tfrac{1}{2}m_{k}\omega_{k}^{2}(r_{k}-\tfrac{g_{k}x}{m_{k}\omega_{k}})^{2}+\frac{p_{k}^{2}}{2m_{k}}\right).$ (33)
-The time evolution of the reduced density matrix is governed by the following master equation [38]
+The time evolution of the reduced density matrix is governed by the following master equation [47, 48]
 $\frac{\partial\rho(x,x^{\prime},t)}{\partial t}$ $=$ $\Bigg\{\frac{-\hbar}{2im}\left(\frac{\partial^{2}}{\partial x^{2}}-\frac{\partial^{2}}{\partial x^{\prime 2}}\right)-\gamma(x-x^{\prime})$
@@
 The field is inhomogeneous in the x-direction.
-Without going into calculational details, we use this setup to illustrate the usefulness the present decoherence measure in dealing with situations where there both discrete and continuous degrees of freedom.
+Without going into calculational details, we use this setup to illustrate the usefulness of the present decoherence measure in dealing with situations where there are both discrete and continuous degrees of freedom.
 Two different kinds of decoherence effects are expected to happen in this experiment, and at two different time scales.
@@
 $\rho$ $=$ $\begin{pmatrix}\rho_{11}&\rho_{12}e^{-i\phi}\\ \rho_{21}e^{i\phi}&\rho_{22}\\ \end{pmatrix}$, (47)
-where $\rho_{11}$ ($\rho_{22}$) represents the probability of the quanton passing through the upper (lower) path. The decoherence measure can then be calculated using (8) as
+where $\rho_{11}$ ($\rho_{22}$) represents the probability of the quanton passing through the upper (lower) path.
+We would like to emphasize here that although the analysis started with a pure quanton state (44), purity of the initial state is not at all a requirement.
+The quanton may already be in a mixed state before entering the setup, but the expression (47) will still hold.
+The phase difference $\phi$ does not represent the phase coherence in the initial state.
+Rather the phase difference may be introduced in the two paths by the phase shifter shown in Fig.
+3.
+The decoherence measure can then be calculated using (8) as
 ${D}_{e}$ $=$ $2(1-\rho_{11}^{2}-\rho_{22}^{2}-2|\rho_{12}|^{2}).$ (48)
-The problem at hand now is find a way to make measurements in the experiment to evaluate the above decoherence measure.
+The problem at hand now is to find a way to make measurements in the experiment to evaluate the above decoherence measure.
 If one blocks the lower path before it enters BS2, there will be equal intensity at D1 and D2 which will be proportional to $\rho_{11}$.
@@
 In this analysis it is assumed that all intensities are normalized with the total intensity.
+One may mention here that even if the incoming quanton is not fully coherent, it will not affect results, because each quanton is split into a superposition of the two paths by a unitary action of the beam splitter BS1.
+Any fluctuation in the phase of the incoming quanton will not affect superposition of the two paths, as that initial phase will factor out of the superposition state.
+Interference will be affected only by what happens after BS1.
+Nevertheless, the results may be affected by some experimental factors such as detector efficiency, and other factors depending on the specific setup.
 Given that the Mach-Zehnder interferometer with cold neutrons is widely used and there have been significant developments in the field of neutron interferometry [49], the experiment suggested here should be feasible with neutrons, allowing for the measurement of decoherence.
@@
 One just needs to evaluate a trace which is basis independent.
+In our opinion, a strong point of the present measure is that it is easily applicable to a wide variety of situations.
+On the other hand, QFI doesn’t measure decoherence directly, rather it measures how sensitive a quantum state is to changes in a chosen parameter.
+Often decrease of QFI may come, not from decoherence, but from some unitary evolution.
+Without careful modeling, QFI cannot cleanly separate noise from coherent dynamics.
 Some earlier works treat the classicality of a state as quantifying the degree of decoherence.
@@
 We believe the measure of decoherence introduced here will turn out to be useful in quantifying decoherence in different situations, especially in the quantum information related experiments.
+While the entanglement between a system and its environment generally cannot be undone, non-Markovianity can throw up interesting situations, especially in scenarios where the distinction between decoherence and dissipation is blurred [55].
+How this decoherence measure holds up in such situations, needs to be explored further.
 Acknowledgements.
 MSQ acknowledges financial support from the University Grants Commission, India, through the Junior Research Fellowship (NTA reference number: 231610153280).
+Author declarations
+The authors have no conflicts to disclose.